A transformer with a primary of 400 turns is connected across a 120v a-c line. the secondary circuit has a potential difference of 300v. the secondary 0.06a and the primary current is 1.85a. how many turns are in the secondary winding? b). what is the transformer efficiency?


1
Expert's answer
2021-06-17T15:29:16-0400

Gives

"N_1=400"

"N_2=?"

"V_1=120V"

"V_2=300"

"I_1=0.06A"

"I_2=1.85A"

"P_1=I_1V_1"

"P_2=V_2I_2"

We know that


Potential proportional to number of turns

"\frac{V_2}{V_1}=\frac{N_2}{N_1}"

"N_2=\frac{V_2}{V_1}\times N_1"

Put value

"N_2=\frac{300}{120}\times400=1000"

"P_1=120\times0.06=7.2W"

"P_2={300}\times1.85=555W"

efficiency="\frac{P_2}{P_1}=\frac{555}{7.2}=77.08"